### v3

Abstract
The advent of the James Webb Space Telescope has revealed a wealth of new galaxies just a few hundred Myr after the Big Bang. Some of these galaxies exhibit unusual N/O ratios that are difficult to explain with stellar populations today. While Wolf-Rayet stars in multiple-burst populations, very massive or rapidly-rotating primordial stars, general relativistic explosions of metal-enriched supermassive stars, or the precursors of globular clusters can in principle account for the nitrogen excess in the galaxies GN-z11 and CEERS 1019, no known stars or supernovae can explain the far higher N/O ratio of 0.46 in GS 3073 at redshift $z=$ 5.55. Here we show that the extreme nitrogen abundances in GS 3073 can be produced by 1000 - 10,000 $M_{\odot}$Ā primordial (Pop III) stars. We find that these are the only candidates that can account for its large N/O ratios and its C/O and Ne/O ratios. GS 3073 is thus the first conclusive evidence in the fossil abundance record of the existence of supermassive Pop III stars at cosmic Dawn.
early universe — dark ages, reionization, first stars — galaxies: formation — galaxies: high-redshift
1 Introduction
Recent surveys with the James Webb Space Telescope (JWST; Finkelstein etĀ al., 2024; Eisenstein etĀ al., 2023; Bezanson etĀ al., 2024) have revealed the existence of high-redshift galaxies with large N/O abundance ratios that are difficult to explain with populations of stars that are known today. Multiple-bursts of massive star formation (Kobayashi & Ferrara, 2024; Rizzuti etĀ al., 2025; McClymont etĀ al., 2025), very massive or rapidly-rotating primordial stars (Nandal etĀ al., 2024b; Tsiatsiou etĀ al., 2024; Nandal etĀ al., 2024c), general relativistic explosions of metal-enriched supermassive stars (SMSs; Nagele & Umeda, 2023)), or the precursors of globular clusters (Cameron etĀ al., 2023; Marques-Chaves etĀ al., 2024) can in principle account for the supersolar nitrogen to oxygen ratios in the galaxies GN-z11 (Bunker etĀ al., 2023; Senchyna etĀ al., 2024) and CEERS 1019 (Larson etĀ al., 2023). However, no known population of stars or supernovae can explain the far higher N/O ratio of 0.46 recently found in GS 3073 at redshift $z=$ 5.55 (Ji etĀ al., 2024).
[…]
2 Numerical Method
We modeled the evolution of 1000 - 10,000 $M_{\odot}$Ā Pop III stars in 1000 $M_{\odot}$Ā increments with the Geneva Stellar Evolution Code (GENEC; Eggenberger etĀ al., 2008).
2.1 GENEC
GENEC is a 1D Lagrangian stellar evolution code that uses the Henyey method (Henyey etĀ al., 1965) to solve the equations of stellar structure. Energy generation and nucleosynthesis is calculated with a 38-isotope reaction network and we use the Helmholz equation of state (EOS; Timmes & Swesty, 2000), which includes contributions from degenerate and non-degenerate relativistic and non-relativistic electrons, electron-positron pair production, and radiation. Internal opacities are derived from the OPAL tables (Rogers & Nayfonov, 2002) and convection is approximated by mixing-length theory.
We initialize the stars as 2 $M_{\odot}$Ā hydrostatic $n\sim 3/2$ polytropes with uniform entropy profiles. They have a homogeneous chemical composition with $X=0.7516$, $Y=0.2484$, and $Z=0$ (Ekstrƶm etĀ al., 2012; Nandal etĀ al., 2024a) and a deuterium mass fraction $X_{2}=5\times 10^{-5}$ (HaemmerlĆ© etĀ al., 2018b; Nandal etĀ al., 2023). They are initially evolved with accretion rates above $2\times 10^{-2}$ $M_{\odot}$Ā yr-1, the critical rate above which the protostars grow as cool, red supergiants (Nandal etĀ al., 2023; Herrington etĀ al., 2023). These rates are later increased as the stars approach the Hayashi limit for the first time to ensure that they reach their target masses by the end of the pre-main sequence or at the onset of core hydrogen burning, after which accretion is terminated. We assume cold accretion, in which the entropy of the accreting material is matched to that of the surface of the star (HaemmerlĆ© etĀ al., 2018b) and is assumed to be radiated away, as in geometrically thin, cold disks (HaemmerlĆ© etĀ al., 2013, 2016). Any excess entropy in the infalling matter is radiated away before reaching the surface of the star (Palla & Stahler, 1992; Hosokawa etĀ al., 2010).
As Pop III stars form from pristine gas, we exclude line-driven winds from our models (Vink etĀ al., 2001). We evolve all ten stars to the end of core oxygen burning and a few to the end of silicon burning. The stars are partitioned into up to 420 zones in mass and are adaptively rezoned over their evolution. The stars were also evolved with adaptive time steps, and a moderate convective overshoot of 0.100 in the pre-main sequence and core hydrogen burning stages but not at later stages. Rotation is not included in these models (HaemmerlĆ© etĀ al., 2018a), but see Nandal etĀ al. (2025) for initial attempts to include rotation in SMS evolution. We apply the first-order post-Newtonian Tolman-Oppenheimer-Volkoff (TOV) correction to the equation of hydrostatic equilibrium throughout the evolution of the stars to accommodate GR effects by modifying Newton’s gravitational constant, $G$ (Fuller etĀ al., 1986; Woods etĀ al., 2017; HaemmerlĆ© etĀ al., 2018b):
$G_{\rm rel}=G\left(1+\frac{P}{\rho c^{2}}+\frac{2GM_{r}}{rc^{2}}+\frac{4\pi Pr^{3}}{M_{r}c^{2}}\right),$ (1)
where $P$ is the pressure, $\rho$ the mass density, $c$ the speed of light and $M_{r}$ is the mass enclosed by radius $r$. We have since confirmed by linear stability analysis that none of the stars in our study encounter the GR instability at any point in their evolution (Saio et al., 2024).
3 Elemental Yields
We show HR and Kippenhahn diagrams for the 7000 $M_{\odot}$Ā star as a fiducial case in FigureĀ 1 and its abundance profiles during He burning and at the end of Si burning in FigureĀ 2. The thin radiative layer between the core and the convective H burning shell in panel (h) of FigureĀ 1 is permeable to products of nuclear burning, which pass into the H shell and fuse with hydrogen or are dispersed throughout the star. 12C from He burning reaches the H shell at 1.22 Myr and fuses into 13N and then 14N. This channel creates the N excess consistent with that observed in GS 3073 because it continues throughout helium burning, the second longest evolutionary stage of the star. N is soon dispersed throughout the star by the numerous convective zones above the H layer, becoming the third most abundant element at the surface as shown in FigureĀ 2. 16O due to alpha captures by 12C in the core, which can also create 14N in the H shell, is also mixed throughout the star, becoming the fourth most abundant element at the surface.
14N and 16O synthesis via the CNO cycle depletes 12C over time. Although it is the fourth most abundant element at the start of helium burning it falls to fifth place below 16O halfway through burning as shown in FigureĀ 2. 14N produced in the H shell also returns to the convective core and forms 22Ne after the onset of He burning. This 22Ne is then mixed throughout the star, becoming the sixth most abundant element at the surface. 16O in the core from the CNO cycle can capture an alpha and form 20Ne, but longer 16O formation times result in less 20Ne being produced than 22Ne. Nuclear ash created in the core after the end of central helium burning never reaches the outer layers of the star because convection timescales are longer than later stages of nuclear burning, which last just a few days. Heavier elements such as 26Mg, 28Si, 44Ti, 48Cr and 56Ni thus remain trapped in the core, as shown in FigureĀ 2. Final yields for all 10 stars and mass losses (as explained below) are shown in Table 1 in the Appendix.
Figure 1: HR (left) and Kippenhahn diagrams (right) for the 7000 $M_{\odot}$Ā star. The entire evolution of the star is shown in (a) and (b), its pre-MS evolution is shown in (c) and (d), its core H burning phase is in (e) and (f), and core He burning to the end of Si burning are in (g) and (h). The black and white lines in the Kippenhahn diagrams are isomass and isothermal lines, respectively, and the red and green dots along the track on the lower left in (a) are the zero-age main sequence (ZAMS) positions of 2 - 120 $M_{\odot}$Ā and 1000 - 10,000 $M_{\odot}$Ā Pop III stars for comparison.
Figure 2: Top: abundance profiles of the 7000 $M_{\odot}$Ā star at the beginning (left) and middle (right) of core He burning. Bottom: abundance profiles of the star at the end of Si burning. Purple and white regions are convective and radiative zones, respectively.
3.1 Abundance Ratios / Dilution Factors
We calculate N/O, C/O, and Ne/O number ratios for comparison to those observed for GS 3073 with
$\log\left(\frac{\frac{M_{X}}{A_{X}}}{\frac{M_{Y}}{A_{Y}}}\right)=\log\left(\frac{M_{X}}{M_{Y}}\right)+\log\left(\frac{A_{Y}}{A_{X}}\right),$ (2)
where the $M$ and $A$ are the masses (in $M_{\odot}$) and mass numbers of the elements, respectively. They are plotted versus mass coordinate for the 7000 $M_{\odot}$Ā star at the end of Si burning in FigureĀ 3, in which the $M$ used in EquationĀ 2 is the total mass of the given element lying above the mass coordinate. As discussed in the Appendix, mass loss from these stars is likely but not well quantified, so we treat it as a free parameter and consider three fiducial values: 10% and 50% of the mass of the star and all the mass above its CO core. Thus, in FigureĀ 3 N/O == 1.175, C/O == -0.48 and O/H ++ 12 is 8.65 if 10% of the mass is ejected. The number ratios rise as the mass coordinate falls to 2050 $M_{\odot}$, the CO boundary, after which they diverge in the core because they depend on the evolutionary stage at which the element was created.
We parametrize enrichment of the galaxy by the star with a dilution factor defined to be the mass of the gas contaminated by ejecta from the star divided by the mass of the ejecta. We show N/O, C/O and Ne/O ratios measured against the O/H ++ 12 ratio for all 10 stars as a function of dilution factor from 1 - 1000 for 10% mass loss, 50% mass loss, and the loss of all mass above the CO core in the left panels of FigureĀ 4. He/H ratios, a prominent observational feature of massive stars, and 12C/13C ratios for all the stars and mass losses are discussed in the Appendix. The green star marks the observed abundance ratio for GS 3073 derived from rest-frame UV emission lines (Ji etĀ al., 2024). The yellow and brown stars are corresponding ratios for CEERS 1019 and GN-z11 (Marques-Chaves etĀ al., 2024; Cameron etĀ al., 2023). Illumination by a Type-1 active galactic nucleus (AGN; Ji etĀ al., 2024) in GS 3073, and its lower redshift, allow these ratios to be tightly constrained.
Figure 3: Element number ratios in the star at the end of Si burning. The masses used to calculate the ratios at a mass coordinate are cumulative, from the stellar atmosphere on the right to the depth corresponding to the given mass loss on the left.
Figure 4: N/O, C/O and Ne/O number ratios for 1000 - 10,000 $M_{\odot}$Ā Pop III stars. Top: 10% mass loss. Center: 50% mass loss. Bottom: loss of all the mass above the CO core. The green, yellow and brown stars indicate corresponding ratios for GS 3073, CEERS 1019 and GN-z11, respectively. The dilution factor is the mass of the gas contaminated by outflows from the star divided by the mass of the outflow.
3.2 N/O Ratios
For mass losses of 10% our stars produce N/O ratios higher than that of GS 3073 except for 9000 $M_{\odot}$, which has an N/O ratio of 0.39 at dilution factors of 20 - 30 that falls within the error bars for GS 3073. The second closest model is the 8000 $M_{\odot}$Ā star, with an N/O ratio of 0.60 at dilution factors of 5 - 15. For mass losses of 50%, 8 of the 10 stars produce N/O ratios above those of GS 3073 except for the 9000 and 10,000 $M_{\odot}$Ā models. The N/O ratios of the 8000 and 9000 $M_{\odot}$Ā stars, 0.52 and 0.41, lie within the error bars for GS 3073 at dilution factors of 24 - 46. If the star sheds all its mass above the CO core only the 3000, 5000 and 7000 $M_{\odot}$Ā stars produce N/O ratios greater than those of GS 3073. The others still exhibit relatively high N/O ratios, such as 0.1 for the 9000 $M_{\odot}$Ā model. The 1000 and 3000 $M_{\odot}$Ā stars closely match the observed abundances, with N/O ratios of 0.38 and 0.48 for dilution factors of 12 - 38 and 31 - 56, respectively.
3.3 C/O Ratios
As shown in the upper right panels of FigureĀ 4, four stars with 10% mass loss produce an overabundance of C to O compared to those of GS 3073. The 2000, 5000 and 7000 $M_{\odot}$Ā stars have C/O ratios that are closest to the observed values, -0.45, -0.44 and -0.14, respectively at dilution factors of 5 - 15 at 2000 $M_{\odot}$Ā but only at a dilution factor of 1 for 5000 and 7000 $M_{\odot}$. For the 50% case, three models produce an overabundance of C/O with respect to GS 3073. The 7000 $M_{\odot}$Ā model most closely matches GS 3073 with a C/O ratio of -0.30 for dilution factors of 1 - 8. For the loss of all mass above the CO core, only 2 stars produce overabundances of C/O, with the 7000 $M_{\odot}$Ā model again most closely matching GS 3073 with a value of -0.28 over dilution factors of 1 - 8.
3.4 Ne/O Ratios
As shown in the lower right panels of FigureĀ 4, the 10,000 $M_{\odot}$Ā star with 10% mass loss is the best match to GS 3073, with Ne/O == -0.71 at a dilution factor of $\sim$ 20. Only one model here produces an overabundance of Ne/O relative to GS 3073. For 50% mass loss the 9000 $M_{\odot}$Ā model is the best match, with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Two of the 10 models produce an overabundance of Ne/O. Finally, for the CO core case two stars match GS 3073, with the 2000 $M_{\odot}$Ā model coming closest with Ne/O == -0.77 at a dilution factor of $\sim$ 20. Four models here produce an overabundance of Ne/O.
3.5 Stars Below 1000 $M_{\odot}$Ā or Above 10,000 $M_{\odot}$
We find that N/O ratios for GS 3073 cannot be explained by Pop III SMSs above 10,000 $M_{\odot}$Ā or below 1000 $M_{\odot}$. 900 and 950 $M_{\odot}$Ā stars evolved until the end of core O burning do not produce N/O ratios greater than -0.16, regardless of mass loss, and 15,000 $M_{\odot}$Ā and 21,000 $M_{\odot}$Ā stars evolved to the end of core C and He burning produce an N/O of at most 0.19, still well below the observed 0.46. The 21,000 $M_{\odot}$Ā star also has a much lower O/H ++12 ratio of 3 - 6, effectively ruling it out as a candidate. These stars strongly suggest the presence of an upper and lower limit in SMS mass beyond which the GS 3073 N/O ratio cannot be produced.
4 Discussion and Conclusion
Figure 5: Abundance ratios of 10 SMSs with contributions by the galactic stars for the case of 10% mass loss. If the chemically mature stellar population contributes C, N, O, and Ne mass fractions $\chi_{\rm C}=$ $2.8\times 10^{-4}$, $\chi_{\rm N}=$ $8\times 10^{-5}$, $\chi_{\rm O}=$ $1.2\times 10^{-3}$ and $\chi_{\rm Ne}=$ $9.4\times 10^{-4}$ then all 10 models can reproduce the observed near vertical transition in N/O ratio while also explaining the C/O ratios (Ji etĀ al., 2024).
No single SMS would be expected to match all the abundance ratios of GS 3073 because other generations of stars also contributed to them. Their contribution decreased N/O ratios and elevated C/O ratios over time because mass loss and SNe produced much more C and O than N (Ekström et al., 2012; Murphy et al., 2021). SMSs with N/O ratios greater than those in GS 3073 and C/O ratios less than those in GS 3073 can thus still be the origin of its N/O excess and C/O fractions. We can estimate the contribution to N/O, C/O and Ne/O ratios in GS 3073 from its less massive stars by comparison to surface abundances of halo and disk stars in the MW at similar metallicities (Prantzos et al., 2018). From Figures 4 and 8 of Ji et al. (2024) we find N and O mass fractions for this population to be 8.0e-5 and 1.2e-3, respectively. Assuming [Fe/H] $\sim$ -0.6 in GS 3073 (Ji et al., 2024), we then obtain C/Fe and Ne/Fe ratios and then C and Ne abundances of 2.8e-4 and 9.4e-4, respectively. We then mix ejecta from the SMS with the ISM for a range of dilution factors.
As shown in FigureĀ 5, when the SMS and stellar contributions to N/O, C/O, and Ne/O are taken together, we obtain good agreement with those of GS 3073 with a 6000 $M_{\odot}$Ā star with 10% mass loss in this scenario. Mixing of the SMS and ISM components produces a nearly vertical shift in N/O versus O/H from its floor of -1.1 to 0.46, suggesting the significant increase in N abundance without a substantial change in O abundance shown in Figure 5 of Ji etĀ al. (2024). This test also shows that the SMSs must be Pop III because no other stars would keep O/H constant. Ours are also the only known stars that can produce the Ne/O ratio in GS 3073. While GR supernovae of metal-enriched SMSs can explain the smaller N excess in GN-z11 (Nagele & Umeda, 2023), they cannot account for its Ne abundances. GR SNe also blow most of the gas from their host galaxies, greatly delaying subsequent star formation (Whalen etĀ al., 2013a, b, d; Johnson etĀ al., 2013; Chen etĀ al., 2014a), so it is not clear if GS 3073 could have reached its stellar mass, 109.5 $M_{\odot}$, by $z=$ 5.55 if a GR SN had occurred early in its history. Furthermore, such an event cannot explain the presence of a BH in GS 3073 because it would not have created one, whereas an SMS can account for both the nitrogen excess in GS 3073 and its AGN. Not only do several of our stars match Ne/O ratios for GS 3073, some produce an overabundance of Ne that would be diluted down to the observed values by other populations of stars.
Gas at lower densities in GS 3073, as traced by optical lines, has a much lower N/O == -1.1 than gas at high densities, which is traced by UV lines (Ji etĀ al., 2024). Stratified N/O ratios are consistent with our SMS scenario, as cosmological simulations indicate that SMSs form in dense gas in high-redshift halos. Whether the SMS formed in a satellite halo and later merged with GS 3073 in which a mature population of stars had already formed (Regan etĀ al., 2017) or it formed in a pristine pocket of gas in GS 3073 (Liu & Bromm, 2020; Venditti etĀ al., 2023), there would have been the near vertical rise in N/O versus O/H. Because the true distribution of low-metallicity and high-metallicity gas in GS 3073 is not known, it is not yet possible to determine precisely which enrichment scenario happened, only that massive nitrogen production by SMSs must have occurred.
The unprecedented supersolar N/O ratio and C/O and Ne/O ratios in GS 3073 are the first conclusive evidence in the fossil abundance record of the existence of supermassive Pop III stars at the end of the cosmic Dark Ages. The N/O overabundances of some of our stars also predict that galaxies with even higher N excesses could exist at high redshifts if they are at earlier stages of star formation. These objects could be found in future surveys by the James Webb Space Telescope and the next generation of ground-based observatories in the decade to come.
DN was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 833925, project STAREX) and by the Virginia Initiative of Theoretical Astronomy (VITA) Fellowship. AH acknowledges support by the Australian Research Council (ARC) through grants DP240101786 and DP240103174. MAL was funded by UAEU via UPAR grant No. 12S111.

### v4

@@
 Abstract
-The advent of the James Webb Space Telescope has revealed a wealth of new galaxies just a few hundred Myr after the Big Bang.
-Some of these galaxies exhibit unusual N/O ratios that are difficult to explain with stellar populations today.
+The advent of the James Webb Space Telescope has revealed a wealth of new galaxies just a few hundred Myr after the Big Bang, a few of which exhibit unusual N/O ratios that are difficult to explain with stellar populations today.
 While Wolf-Rayet stars in multiple-burst populations, very massive or rapidly-rotating primordial stars, general relativistic explosions of metal-enriched supermassive stars, or the precursors of globular clusters can in principle account for the nitrogen excess in the galaxies GN-z11 and CEERS 1019, no known stars or supernovae can explain the far higher N/O ratio of 0.46 in GS 3073 at redshift $z=$ 5.55.
-Here we show that the extreme nitrogen abundances in GS 3073 can be produced by 1000 - 10,000 $M_{\odot}$Ā primordial (Pop III) stars.
+Here we show that the extreme nitrogen abundances in GS 3073 can be produced by 1000 - 10,000 $M_{\odot}$ primordial (Pop III) stars.
 We find that these are the only candidates that can account for its large N/O ratios and its C/O and Ne/O ratios.
@@
 1 Introduction
-Recent surveys with the James Webb Space Telescope (JWST; Finkelstein etĀ al., 2024; Eisenstein etĀ al., 2023; Bezanson etĀ al., 2024) have revealed the existence of high-redshift galaxies with large N/O abundance ratios that are difficult to explain with populations of stars that are known today.
-Multiple-bursts of massive star formation (Kobayashi & Ferrara, 2024; Rizzuti etĀ al., 2025; McClymont etĀ al., 2025), very massive or rapidly-rotating primordial stars (Nandal etĀ al., 2024b; Tsiatsiou etĀ al., 2024; Nandal etĀ al., 2024c), general relativistic explosions of metal-enriched supermassive stars (SMSs; Nagele & Umeda, 2023)), or the precursors of globular clusters (Cameron etĀ al., 2023; Marques-Chaves etĀ al., 2024) can in principle account for the supersolar nitrogen to oxygen ratios in the galaxies GN-z11 (Bunker etĀ al., 2023; Senchyna etĀ al., 2024) and CEERS 1019 (Larson etĀ al., 2023).
-However, no known population of stars or supernovae can explain the far higher N/O ratio of 0.46 recently found in GS 3073 at redshift $z=$ 5.55 (Ji etĀ al., 2024).
+Recent surveys with the James Webb Space Telescope (JWST; Finkelstein et al., 2024; Eisenstein et al., 2023; Bezanson et al., 2024) have revealed the existence of high-redshift galaxies with large N/O abundance ratios that are difficult to explain with populations of stars that are known today.
+Multiple-bursts of massive star formation (Kobayashi & Ferrara, 2024; Rizzuti et al., 2025; McClymont et al., 2025), very massive or rapidly-rotating primordial stars (Nandal et al., 2024b; Tsiatsiou et al., 2024; Nandal et al., 2024c), general relativistic explosions of metal-enriched supermassive stars (SMSs; Nagele & Umeda, 2023)), or the precursors of globular clusters (Cameron et al., 2023; Marques-Chaves et al., 2024) can in principle account for the supersolar nitrogen to oxygen ratios in the galaxies GN-z11 (Bunker et al., 2023; Senchyna et al., 2024) and CEERS 1019 (Larson et al., 2023).
+However, no known population of stars or supernovae can explain the far higher N/O ratio of 0.46 recently found in GS 3073 at redshift $z=$ 5.55 (Ji et al., 2024).
 […]
 2 Numerical Method
-We modeled the evolution of 1000 - 10,000 $M_{\odot}$Ā Pop III stars in 1000 $M_{\odot}$Ā increments with the Geneva Stellar Evolution Code (GENEC; Eggenberger etĀ al., 2008).
+We modeled the evolution of 1000 - 10,000 $M_{\odot}$ Pop III stars in 1000 $M_{\odot}$ increments with the Geneva Stellar Evolution Code (GENEC; Eggenberger et al., 2008).
 2.1 GENEC
-GENEC is a 1D Lagrangian stellar evolution code that uses the Henyey method (Henyey etĀ al., 1965) to solve the equations of stellar structure.
+GENEC is a 1D Lagrangian stellar evolution code that uses the Henyey method (Henyey et al., 1965) to solve the equations of stellar structure.
 Energy generation and nucleosynthesis is calculated with a 38-isotope reaction network and we use the Helmholz equation of state (EOS; Timmes & Swesty, 2000), which includes contributions from degenerate and non-degenerate relativistic and non-relativistic electrons, electron-positron pair production, and radiation.
 Internal opacities are derived from the OPAL tables (Rogers & Nayfonov, 2002) and convection is approximated by mixing-length theory.
-We initialize the stars as 2 $M_{\odot}$Ā hydrostatic $n\sim 3/2$ polytropes with uniform entropy profiles.
-They have a homogeneous chemical composition with $X=0.7516$, $Y=0.2484$, and $Z=0$ (Ekstrƶm etĀ al., 2012; Nandal etĀ al., 2024a) and a deuterium mass fraction $X_{2}=5\times 10^{-5}$ (HaemmerlĆ© etĀ al., 2018b; Nandal etĀ al., 2023).
-They are initially evolved with accretion rates above $2\times 10^{-2}$ $M_{\odot}$Ā yr-1, the critical rate above which the protostars grow as cool, red supergiants (Nandal etĀ al., 2023; Herrington etĀ al., 2023).
+We initialize the stars as 2 $M_{\odot}$ hydrostatic $n\sim 3/2$ polytropes with uniform entropy profiles.
+They have a homogeneous chemical composition with $X=0.7516$, $Y=0.2484$, and $Z=0$ (Ekström et al., 2012; Nandal et al., 2024a) and a deuterium mass fraction $X_{2}=5\times 10^{-5}$ (Haemmerlé et al., 2018b; Nandal et al., 2023).
+They are initially evolved with accretion rates above $2\times 10^{-2}$ $M_{\odot}$ yr-1, the critical rate above which the protostars grow as cool, red supergiants (Hosokawa et al., 2012, 2013; Nandal et al., 2023; Herrington et al., 2023).
 These rates are later increased as the stars approach the Hayashi limit for the first time to ensure that they reach their target masses by the end of the pre-main sequence or at the onset of core hydrogen burning, after which accretion is terminated.
-We assume cold accretion, in which the entropy of the accreting material is matched to that of the surface of the star (HaemmerlĆ© etĀ al., 2018b) and is assumed to be radiated away, as in geometrically thin, cold disks (HaemmerlĆ© etĀ al., 2013, 2016).
-Any excess entropy in the infalling matter is radiated away before reaching the surface of the star (Palla & Stahler, 1992; Hosokawa etĀ al., 2010).
-As Pop III stars form from pristine gas, we exclude line-driven winds from our models (Vink etĀ al., 2001).
+We assume cold accretion, in which the entropy of the accreting material is matched to that of the surface of the star (Haemmerlé et al., 2018b) and is assumed to be radiated away, as in geometrically thin, cold disks (Haemmerlé et al., 2013, 2016).
+Any excess entropy in the infalling matter is radiated away before reaching the surface of the star (Palla & Stahler, 1992; Hosokawa et al., 2010).
+For fixed final mass and fixed ejection rule, the integrated ejecta are insensitive to whether pre–MS accretion is variable or constant.
+By the end of core–He burning the differences are $\leq 10^{-2}$ in mass fraction for N, O, and Ne, and by core–Si burning they are $\leq 10^{-4}$; the only systematic change is the total age.
+As Pop III stars form from pristine gas, we exclude line-driven winds from our models (Vink et al., 2001).
 We evolve all ten stars to the end of core oxygen burning and a few to the end of silicon burning.
@@
 The stars were also evolved with adaptive time steps, and a moderate convective overshoot of 0.100 in the pre-main sequence and core hydrogen burning stages but not at later stages.
-Rotation is not included in these models (HaemmerlĆ© etĀ al., 2018a), but see Nandal etĀ al.
+Rotation is not included in these models (Haemmerlé et al., 2018a), but see Nandal et al.
 (2025) for initial attempts to include rotation in SMS evolution.
-We apply the first-order post-Newtonian Tolman-Oppenheimer-Volkoff (TOV) correction to the equation of hydrostatic equilibrium throughout the evolution of the stars to accommodate GR effects by modifying Newton’s gravitational constant, $G$ (Fuller etĀ al., 1986; Woods etĀ al., 2017; HaemmerlĆ© etĀ al., 2018b):
+We apply the first-order post-Newtonian Tolman-Oppenheimer-Volkoff (TOV) correction to the equation of hydrostatic equilibrium throughout the evolution of the stars to accommodate GR effects by modifying Newton’s gravitational constant, $G$ (Fuller et al., 1986; Woods et al., 2017; Haemmerlé et al., 2018b):
 $G_{\rm rel}=G\left(1+\frac{P}{\rho c^{2}}+\frac{2GM_{r}}{rc^{2}}+\frac{4\pi Pr^{3}}{M_{r}c^{2}}\right),$ (1)
@@
 We have since confirmed by linear stability analysis that none of the stars in our study encounter the GR instability at any point in their evolution (Saio et al., 2024).
+In contrast to Nandal et al.
+(2024b), we follow each model at least to core–O burning and for a subset to core–Si burning, which fixes the O and Ne stratification outside the CO core and enables robust Ne/O predictions used below.
 3 Elemental Yields
-We show HR and Kippenhahn diagrams for the 7000 $M_{\odot}$Ā star as a fiducial case in FigureĀ 1 and its abundance profiles during He burning and at the end of Si burning in FigureĀ 2.
-The thin radiative layer between the core and the convective H burning shell in panel (h) of FigureĀ 1 is permeable to products of nuclear burning, which pass into the H shell and fuse with hydrogen or are dispersed throughout the star.
+Figure 1: HR (top) and Kippenhahn diagrams (bottom) for the 7000 $M_{\odot}$ star.
+The entire evolution of the star is shown in (a) and (b), its pre-MS evolution is shown in (c) and (d), its core H burning phase is in (e) and (f), and core He burning to the end of Si burning are in (g) and (h).
+The black and white lines in the Kippenhahn diagrams are isomass and isothermal lines, respectively, and the red and green dots along the track on the lower left in (a) are the zero-age main sequence (ZAMS) positions of 2 - 120 $M_{\odot}$ and 1000 - 10,000 $M_{\odot}$ Pop III stars for comparison.
+Figure 2: Abundance profiles of the $7000\,M_{\odot}$ star at three evolutionary stages.
+Top: beginning of core He burning.
+Middle: middle of core He burning.
+Bottom: end of Si burning.
+Purple and yellow regions mark convective and radiative zones, respectively.
+We show HR and Kippenhahn diagrams for the 7000 $M_{\odot}$ star as a fiducial case in Figure 1 and its abundance profiles during He burning and at the end of Si burning in Figure 2.
+The thin radiative layer between the core and the convective H burning shell in panel (h) of Figure 1 is permeable to products of nuclear burning, which pass into the H shell and fuse with hydrogen or are dispersed throughout the star.
 12C from He burning reaches the H shell at 1.22 Myr and fuses into 13N and then 14N.
 This channel creates the N excess consistent with that observed in GS 3073 because it continues throughout helium burning, the second longest evolutionary stage of the star.
-N is soon dispersed throughout the star by the numerous convective zones above the H layer, becoming the third most abundant element at the surface as shown in FigureĀ 2.
+N is soon dispersed throughout the star by the numerous convective zones above the H layer, becoming the third most abundant element at the surface as shown in Figure 2.
 16O due to alpha captures by 12C in the core, which can also create 14N in the H shell, is also mixed throughout the star, becoming the fourth most abundant element at the surface.
 14N and 16O synthesis via the CNO cycle depletes 12C over time.
-Although it is the fourth most abundant element at the start of helium burning it falls to fifth place below 16O halfway through burning as shown in FigureĀ 2.
+Although it is the fourth most abundant element at the start of helium burning it falls to fifth place below 16O halfway through burning as shown in Figure 2.
 14N produced in the H shell also returns to the convective core and forms 22Ne after the onset of He burning.
@@
 Nuclear ash created in the core after the end of central helium burning never reaches the outer layers of the star because convection timescales are longer than later stages of nuclear burning, which last just a few days.
-Heavier elements such as 26Mg, 28Si, 44Ti, 48Cr and 56Ni thus remain trapped in the core, as shown in FigureĀ 2.
+Heavier elements such as 26Mg, 28Si, 44Ti, 48Cr and 56Ni thus remain trapped in the core, as shown in Figure 2.
 Final yields for all 10 stars and mass losses (as explained below) are shown in Table 1 in the Appendix.
-Figure 1: HR (left) and Kippenhahn diagrams (right) for the 7000 $M_{\odot}$Ā star.
-The entire evolution of the star is shown in (a) and (b), its pre-MS evolution is shown in (c) and (d), its core H burning phase is in (e) and (f), and core He burning to the end of Si burning are in (g) and (h).
-The black and white lines in the Kippenhahn diagrams are isomass and isothermal lines, respectively, and the red and green dots along the track on the lower left in (a) are the zero-age main sequence (ZAMS) positions of 2 - 120 $M_{\odot}$Ā and 1000 - 10,000 $M_{\odot}$Ā Pop III stars for comparison.
-Figure 2: Top: abundance profiles of the 7000 $M_{\odot}$Ā star at the beginning (left) and middle (right) of core He burning.
-Bottom: abundance profiles of the star at the end of Si burning.
-Purple and white regions are convective and radiative zones, respectively.
 3.1 Abundance Ratios / Dilution Factors
@@
 where the $M$ and $A$ are the masses (in $M_{\odot}$) and mass numbers of the elements, respectively.
-They are plotted versus mass coordinate for the 7000 $M_{\odot}$Ā star at the end of Si burning in FigureĀ 3, in which the $M$ used in EquationĀ 2 is the total mass of the given element lying above the mass coordinate.
+They are plotted versus mass coordinate for the 7000 $M_{\odot}$ star at the end of Si burning in Figure 3, in which the $M$ used in Equation 2 is the total mass of the given element lying above the mass coordinate.
 As discussed in the Appendix, mass loss from these stars is likely but not well quantified, so we treat it as a free parameter and consider three fiducial values: 10% and 50% of the mass of the star and all the mass above its CO core.
-Thus, in FigureĀ 3 N/O == 1.175, C/O == -0.48 and O/H ++ 12 is 8.65 if 10% of the mass is ejected.
+Thus, in Figure 3 N/O == 1.175, C/O == -0.48 and O/H ++ 12 is 8.65 if 10% of the mass is ejected.
 The number ratios rise as the mass coordinate falls to 2050 $M_{\odot}$, the CO boundary, after which they diverge in the core because they depend on the evolutionary stage at which the element was created.
 We parametrize enrichment of the galaxy by the star with a dilution factor defined to be the mass of the gas contaminated by ejecta from the star divided by the mass of the ejecta.
-We show N/O, C/O and Ne/O ratios measured against the O/H ++ 12 ratio for all 10 stars as a function of dilution factor from 1 - 1000 for 10% mass loss, 50% mass loss, and the loss of all mass above the CO core in the left panels of FigureĀ 4.
+We show N/O, C/O and Ne/O ratios measured against the O/H ++ 12 ratio for all 10 stars as a function of dilution factor from 1 - 1000 for 10% mass loss, 50% mass loss, and the loss of all mass above the CO core in the left panels of Figure 4.
 He/H ratios, a prominent observational feature of massive stars, and 12C/13C ratios for all the stars and mass losses are discussed in the Appendix.
-The green star marks the observed abundance ratio for GS 3073 derived from rest-frame UV emission lines (Ji etĀ al., 2024).
-The yellow and brown stars are corresponding ratios for CEERS 1019 and GN-z11 (Marques-Chaves etĀ al., 2024; Cameron etĀ al., 2023).
-Illumination by a Type-1 active galactic nucleus (AGN; Ji etĀ al., 2024) in GS 3073, and its lower redshift, allow these ratios to be tightly constrained.
+The green star marks the observed abundance ratio for GS 3073 derived from rest-frame UV emission lines (Ji et al., 2024).
+The yellow and brown stars are corresponding ratios for CEERS 1019 and GN-z11 (Marques-Chaves et al., 2024; Cameron et al., 2023).
+Illumination by a Type-1 active galactic nucleus (AGN; Ji et al., 2024) in GS 3073, and its lower redshift, allow these ratios to be tightly constrained.
+Nitrogen enrichment follows a robust sequence in our supermassive models: 12C produced in the convective He core permeates the thin radiative buffer into the convective H–burning shell, where the CNO cycle converts it to 14N; intermediate convective zones above the shell then disperse 14N through the envelope (see Figs.
+1 and 2).
+The appearance of these zones is controlled by the standard convective instability criterion, $\nabla_{\rm rad}>\nabla_{\rm ad}$, with the radiative gradient
+$\nabla_{\rm rad}\;=\;\frac{3\,\kappa\,L\,P}{16\,\pi\,a\,c\,G\,m\,T^{4}}\,$,
+so that near–Eddington luminosities in SMS envelopes favor extended convection without requiring any ad hoc mixing.
+After accretion ceases, the structure relaxes on a Kelvin–Helmholtz timescale $\tau_{\rm KH}\!=\!GM^{2}/(RL)$ that is short compared to the duration of core-He burning; as a result the convective topology and the 12C supply converge across $10^{3}$$10^{4}$ M.
+This produces high N/O with only weak stellar mass dependence in the fixed fraction mass-loss cases (Top and Center panels of Fig.
+4).
+By contrast, in the “above CO core” prescription a small change in the mass cut across the steep N rich/O poor layer just outside the CO core can yield large local excursions in N/O, reflecting a dependence on the composition gradient, rather than on a progenitor mass trend.
 Figure 3: Element number ratios in the star at the end of Si burning.
 The masses used to calculate the ratios at a mass coordinate are cumulative, from the stellar atmosphere on the right to the depth corresponding to the given mass loss on the left.
-Figure 4: N/O, C/O and Ne/O number ratios for 1000 - 10,000 $M_{\odot}$Ā Pop III stars.
+Figure 4: N/O, C/O and Ne/O number ratios for 5 select models in the mass range of 1000 - 10,000 $M_{\odot}$ Pop III stars.
 Top: 10% mass loss.
@@
 The dilution factor is the mass of the gas contaminated by outflows from the star divided by the mass of the outflow.
+Outliers appear only in the “CO core” case (bottom panel), where the mass cut can intersect the narrow N rich/O poor layer immediately outside the core; the 10% and 50% fixed fraction cases average over this gradient and exhibit much smaller scatter.
 3.2 N/O Ratios
 For mass losses of 10% our stars produce N/O ratios higher than that of GS 3073 except for 9000 $M_{\odot}$, which has an N/O ratio of 0.39 at dilution factors of 20 - 30 that falls within the error bars for GS 3073.
-The second closest model is the 8000 $M_{\odot}$Ā star, with an N/O ratio of 0.60 at dilution factors of 5 - 15.
-For mass losses of 50%, 8 of the 10 stars produce N/O ratios above those of GS 3073 except for the 9000 and 10,000 $M_{\odot}$Ā models.
-The N/O ratios of the 8000 and 9000 $M_{\odot}$Ā stars, 0.52 and 0.41, lie within the error bars for GS 3073 at dilution factors of 24 - 46.
-If the star sheds all its mass above the CO core only the 3000, 5000 and 7000 $M_{\odot}$Ā stars produce N/O ratios greater than those of GS 3073.
-The others still exhibit relatively high N/O ratios, such as 0.1 for the 9000 $M_{\odot}$Ā model.
-The 1000 and 3000 $M_{\odot}$Ā stars closely match the observed abundances, with N/O ratios of 0.38 and 0.48 for dilution factors of 12 - 38 and 31 - 56, respectively.
+The second closest model is the 8000 $M_{\odot}$ star, with an N/O ratio of 0.60 at dilution factors of 5 - 15.
+For mass losses of 50%, 8 of the 10 stars produce N/O ratios above those of GS 3073 except for the 9000 and 10,000 $M_{\odot}$ models.
+The N/O ratios of the 8000 and 9000 $M_{\odot}$ stars, 0.52 and 0.41, lie within the error bars for GS 3073 at dilution factors of 24 - 46.
+If the star sheds all its mass above the CO core only the 3000, 5000 and 7000 $M_{\odot}$ stars produce N/O ratios greater than those of GS 3073.
+The others still exhibit relatively high N/O ratios, such as 0.1 for the 9000 $M_{\odot}$ model.
+The 1000 and 3000 $M_{\odot}$ stars closely match the observed abundances, with N/O ratios of 0.38 and 0.48 for dilution factors of 12 - 38 and 31 - 56, respectively.
+The extended evolution allows us to match the observed $\log(\mathrm{Ne/O})$ of GS 3073 while preserving the near–vertical rise in N/O at roughly fixed O/H (Ji et al., 2024); this diagnostic was unavailable in our earlier work because the models did not reach advanced burning.
 3.3 C/O Ratios
-As shown in the upper right panels of FigureĀ 4, four stars with 10% mass loss produce an overabundance of C to O compared to those of GS 3073.
-The 2000, 5000 and 7000 $M_{\odot}$Ā stars have C/O ratios that are closest to the observed values, -0.45, -0.44 and -0.14, respectively at dilution factors of 5 - 15 at 2000 $M_{\odot}$Ā but only at a dilution factor of 1 for 5000 and 7000 $M_{\odot}$.
+As shown in the upper right panels of Figure 4, four stars with 10% mass loss produce an overabundance of C to O compared to those of GS 3073.
+The 2000, 5000 and 7000 $M_{\odot}$ stars have C/O ratios that are closest to the observed values, -0.45, -0.44 and -0.14, respectively at dilution factors of 5 - 15 at 2000 $M_{\odot}$ but only at a dilution factor of 1 for 5000 and 7000 $M_{\odot}$.
 For the 50% case, three models produce an overabundance of C/O with respect to GS 3073.
-The 7000 $M_{\odot}$Ā model most closely matches GS 3073 with a C/O ratio of -0.30 for dilution factors of 1 - 8.
-For the loss of all mass above the CO core, only 2 stars produce overabundances of C/O, with the 7000 $M_{\odot}$Ā model again most closely matching GS 3073 with a value of -0.28 over dilution factors of 1 - 8.
+The 7000 $M_{\odot}$ model most closely matches GS 3073 with a C/O ratio of -0.30 for dilution factors of 1 - 8.
+For the loss of all mass above the CO core, only 2 stars produce overabundances of C/O, with the 7000 $M_{\odot}$ model again most closely matching GS 3073 with a value of -0.28 over dilution factors of 1 - 8.
 3.4 Ne/O Ratios
-As shown in the lower right panels of FigureĀ 4, the 10,000 $M_{\odot}$Ā star with 10% mass loss is the best match to GS 3073, with Ne/O == -0.71 at a dilution factor of $\sim$ 20.
+As shown in the lower right panels of Figure 4, the 10,000 $M_{\odot}$ star with 10% mass loss is the best match to GS 3073, with Ne/O == -0.71 at a dilution factor of $\sim$ 20.
 Only one model here produces an overabundance of Ne/O relative to GS 3073.
-For 50% mass loss the 9000 $M_{\odot}$Ā model is the best match, with Ne/O == -0.77 at a dilution factor of $\sim$ 20.
+For 50% mass loss the 9000 $M_{\odot}$ model is the best match, with Ne/O == -0.77 at a dilution factor of $\sim$ 20.
 Two of the 10 models produce an overabundance of Ne/O.
-Finally, for the CO core case two stars match GS 3073, with the 2000 $M_{\odot}$Ā model coming closest with Ne/O == -0.77 at a dilution factor of $\sim$ 20.
+Finally, for the CO core case two stars match GS 3073, with the 2000 $M_{\odot}$ model coming closest with Ne/O == -0.77 at a dilution factor of $\sim$ 20.
 Four models here produce an overabundance of Ne/O.
-3.5 Stars Below 1000 $M_{\odot}$Ā or Above 10,000 $M_{\odot}$
-We find that N/O ratios for GS 3073 cannot be explained by Pop III SMSs above 10,000 $M_{\odot}$Ā or below 1000 $M_{\odot}$.
-900 and 950 $M_{\odot}$Ā stars evolved until the end of core O burning do not produce N/O ratios greater than -0.16, regardless of mass loss, and 15,000 $M_{\odot}$Ā and 21,000 $M_{\odot}$Ā stars evolved to the end of core C and He burning produce an N/O of at most 0.19, still well below the observed 0.46.
-The 21,000 $M_{\odot}$Ā star also has a much lower O/H ++12 ratio of 3 - 6, effectively ruling it out as a candidate.
+3.5 Stars Below 1000 $M_{\odot}$ or Above 10,000 $M_{\odot}$
+We find that N/O ratios for GS 3073 cannot be explained by Pop III SMSs above 10,000 $M_{\odot}$ or below 1000 $M_{\odot}$.
+900 and 950 $M_{\odot}$ stars evolved until the end of core O burning do not produce N/O ratios greater than -0.16, regardless of mass loss, and 15,000 $M_{\odot}$ and 21,000 $M_{\odot}$ stars evolved to the end of core C and He burning produce an N/O of at most 0.19, still well below the observed 0.46.
+The 21,000 $M_{\odot}$ star also has a much lower O/H ++12 ratio of 3 - 6, effectively ruling it out as a candidate.
 These stars strongly suggest the presence of an upper and lower limit in SMS mass beyond which the GS 3073 N/O ratio cannot be produced.
+The emergence of this window is also captured by the integrated ejecta shown in Table 1. Defining
+$M_{\rm N}^{\rm ej}$ $=\int_{m_{\rm cut}}^{M}X_{\rm N}(m)\,\mathrm{d}m,$
+$M_{\rm O}^{\rm ej}$ $=\int_{m_{\rm cut}}^{M}X_{\rm O}(m)\,\mathrm{d}m.$
+[…]
 4 Discussion and Conclusion
-Figure 5: Abundance ratios of 10 SMSs with contributions by the galactic stars for the case of 10% mass loss.
-If the chemically mature stellar population contributes C, N, O, and Ne mass fractions $\chi_{\rm C}=$ $2.8\times 10^{-4}$, $\chi_{\rm N}=$ $8\times 10^{-5}$, $\chi_{\rm O}=$ $1.2\times 10^{-3}$ and $\chi_{\rm Ne}=$ $9.4\times 10^{-4}$ then all 10 models can reproduce the observed near vertical transition in N/O ratio while also explaining the C/O ratios (Ji etĀ al., 2024).
+Figure 5: Abundance ratios of 5 SMSs with contributions by the galactic stars for the case of 10% mass loss.
+The green, yellow and brown stars indicate corresponding ratios for GS 3073, CEERS 1019 and GN-z11, respectively.
+If the chemically mature stellar population contributes C, N, O, and Ne mass fractions $\chi_{\rm C}=$ $2.8\times 10^{-4}$, $\chi_{\rm N}=$ $8\times 10^{-5}$, $\chi_{\rm O}=$ $1.2\times 10^{-3}$ and $\chi_{\rm Ne}=$ $9.4\times 10^{-4}$ then all 10 models can reproduce the observed near vertical transition in N/O ratio while also explaining the C/O ratios (Ji et al., 2024).
 No single SMS would be expected to match all the abundance ratios of GS 3073 because other generations of stars also contributed to them.
@@
 We then mix ejecta from the SMS with the ISM for a range of dilution factors.
-As shown in FigureĀ 5, when the SMS and stellar contributions to N/O, C/O, and Ne/O are taken together, we obtain good agreement with those of GS 3073 with a 6000 $M_{\odot}$Ā star with 10% mass loss in this scenario.
-Mixing of the SMS and ISM components produces a nearly vertical shift in N/O versus O/H from its floor of -1.1 to 0.46, suggesting the significant increase in N abundance without a substantial change in O abundance shown in Figure 5 of Ji etĀ al.
+The present grid differs from Nandal et al.
+(2024b) by advancing to core–O (and in some cases core–Si) burning with TOV corrections, by adding Ne/O as a key abundance test, by confronting the tighter abundance set of GS 3073, and by including an exploratory rotating SMS subset to assess sensitivity to rotational mixing during He burning.
+Additionally, we also explain the black hole mass of GS3073 by invoking sub-Eddington accretion onto the seed formed by such SMSs at redshifts up to 20.
+As shown in Figure 5, when the SMS and stellar contributions to N/O, C/O, and Ne/O are taken together, we obtain good agreement with those of GS 3073 with a 6000 $M_{\odot}$ star with 10% mass loss in this scenario.
+Mixing of the SMS and ISM components produces a nearly vertical shift in N/O versus O/H from its floor of -1.1 to 0.46, suggesting the significant increase in N abundance without a substantial change in O abundance shown in Figure 5 of Ji et al.
 (2024).
@@
 While GR supernovae of metal-enriched SMSs can explain the smaller N excess in GN-z11 (Nagele & Umeda, 2023), they cannot account for its Ne abundances.
-GR SNe also blow most of the gas from their host galaxies, greatly delaying subsequent star formation (Whalen etĀ al., 2013a, b, d; Johnson etĀ al., 2013; Chen etĀ al., 2014a), so it is not clear if GS 3073 could have reached its stellar mass, 109.5 $M_{\odot}$, by $z=$ 5.55 if a GR SN had occurred early in its history.
+GR SNe also blow most of the gas from their host galaxies, greatly delaying subsequent star formation (Whalen et al., 2013a, b, d; Johnson et al., 2013; Chen et al., 2014a), so it is not clear if GS 3073 could have reached its stellar mass, 109.5 $M_{\odot}$, by $z=$ 5.55 if a GR SN had occurred early in its history.
 Furthermore, such an event cannot explain the presence of a BH in GS 3073 because it would not have created one, whereas an SMS can account for both the nitrogen excess in GS 3073 and its AGN.
 Not only do several of our stars match Ne/O ratios for GS 3073, some produce an overabundance of Ne that would be diluted down to the observed values by other populations of stars.
-Gas at lower densities in GS 3073, as traced by optical lines, has a much lower N/O == -1.1 than gas at high densities, which is traced by UV lines (Ji etĀ al., 2024).
+Gas at lower densities in GS 3073, as traced by optical lines, has a much lower N/O == -1.1 than gas at high densities, which is traced by UV lines (Ji et al., 2024).
 Stratified N/O ratios are consistent with our SMS scenario, as cosmological simulations indicate that SMSs form in dense gas in high-redshift halos.
-Whether the SMS formed in a satellite halo and later merged with GS 3073 in which a mature population of stars had already formed (Regan etĀ al., 2017) or it formed in a pristine pocket of gas in GS 3073 (Liu & Bromm, 2020; Venditti etĀ al., 2023), there would have been the near vertical rise in N/O versus O/H.
+Whether the SMS formed in a satellite halo and later merged with GS 3073 in which a mature population of stars had already formed (Regan et al., 2017) or it formed in a pristine pocket of gas in GS 3073 (Liu & Bromm, 2020; Venditti et al., 2023), there would have been the near vertical rise in N/O versus O/H.
 Because the true distribution of low-metallicity and high-metallicity gas in GS 3073 is not known, it is not yet possible to determine precisely which enrichment scenario happened, only that massive nitrogen production by SMSs must have occurred.
@@
 These objects could be found in future surveys by the James Webb Space Telescope and the next generation of ground-based observatories in the decade to come.
-DN was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 833925, project STAREX) and by the Virginia Initiative of Theoretical Astronomy (VITA) Fellowship.
+The authors would like to thank the referee for their invaluable comments and feedback on the manuscript.
+DN was supported by the Swiss National Science Fund (SNSF) Postdoctoral Fellowship, grant number: P500-2235464 and by the Virginia Initiative of Theoretical Astronomy (VITA) Fellowship.
 AH acknowledges support by the Australian Research Council (ARC) through grants DP240101786 and DP240103174.